OpenStudy (anonymous):

A crystal ball with a diameter of 2r inches is being packaged for shipment. If the crystal ball is placed inside a circular cylinder with radius r inches and height h inches, how much volume will need to be filled with padding? Assume r = 6, h = 12. (The volume of a sphere with radius r is 4πr3/3, and the volume of a right circular cylinder with radius r and height h is πr 2h.)

Directrix (directrix):

Volume of Cylinder is pi r^2 h V = pi (6)^2 (12) V = 432 pi Volume of sphere = 4/3 pi r^3 V = 4/3 pi (6)^3 V = 288 pi Space for Packaging = 432 pi - 288 pi = 144 pi (Check my work)

OpenStudy (anonymous):

Thank you so much!!!